• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 00:03:13 on March 15, 2006

    Entry number 166705

    keyword=Beam Time Accounting
    Hall A

    03/14/06 1600-2400
    Experiment #: E02-013
    Shift Leader: A. Puckett
    Shift Workers: V. Punjabi

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0.000.200.060.380.150.390.000.361.54
    Beam Available
    but Not in Use
    0.000.000.890.220.790.320.980.563.76
    Beam Not Available
    or Unacceptable
    1.000.800.050.400.060.290.020.082.70
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    1.001.000.120.780.220.680.000.434.23
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.000.000.880.220.780.321.000.573.77
    Hall not
    ready reasons
    Target0.000.000.000.000.000.321.000.571.88
    Data Acq0.000.000.000.000.000.000.000.000.00
    Other0.000.000.880.220.780.000.000.001.88
    HRSL0.000.000.000.000.000.000.000.000.00
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: Between 20:00 and 21:00 beam was not really acceptable. Positions were eventually corrected but there was substantial beam halo as the Compton backgrounds were unacceptably high.

    Data Entry by: A. Puckett